# Unified Multimodal Understanding and Generation Models: Advances, Challenges, and Opportunities

### v3

@@ 2.2 Text-to-Image Model
3 Unified Multimodal Models for Understanding and Generation
3.1 Diffusion Models
Despite the advances brought by models like Dual Diffusion and MMaDA, several limitations remain. First, computational inefficiency is still a major concern. These models rely on lengthy iterative denoising processes across multiple modalities, making them slower than GANs or autoregressive approaches in real-time applications. Second, while cross-modal conditioning enhances alignment, it remains sensitive to noise schedules and latent inconsistency, especially when handling high-entropy visual or textual inputs. In MMaDA, the integration of chain-of-thought reasoning and reinforcement learning adds further overhead and may suffer from reward sparsity or misalignment across modalities during fine-tuning. Lastly, both Dual Diffusion and MMaDA still struggle with fine-grained control over generation and suffer form limited generalization to novel concepts or out-of-distribution prompts, indicating that robust multimodal abstraction and transfer remain open challenges.
3.2 Auto-Regressive Models
2) Semantic Encoding. To overcome the semantic limitations inherent in pixel-based encoders, a growing body of work adopts semantic encoding, where image inputs are processed using pretrained text-aligned vision encoders such as OpenAI-CLIP [22], SigLIP [152], EVA-CLIP [36], or more recent unified tokenizers like UNIT [154], as shown in Fig. 5 (b-2). These models are trained on large-scale image-text pairs with contrastive or regression-based objectives, producing visual embeddings that align closely with language features in a shared semantic space. Such representations enable more effective cross-modal alignment and are particularly beneficial for multimodal understanding and generation.
Several representative models leverage different semantic encoders and architectural designs to support unified multimodal tasks. Emu [155], Emu2 [33], and LaViT [156] all employ EVA-CLIP [36] as their vision encoder. Notably, Emu [155] introduces the initial architecture combining a frozen EVA-CLIP encoder, a large language model, and a diffusion decoder to unify VQA, image captioning, and image generation. Emu2 [33] builds upon Emu [155] by proposing a simplified and scalable modeling framework for unified multimodal pretraining. It scales the MLLM model up to 37B parameters, significantly enhancing both understanding and generation capabilities. LaViT [156] introduces a dynamic visual tokenization mechanism built on top of EVA-CLIP. It employs a selector and merger module to adaptively select visual tokens from image embeddings based on content complexity. This process dynamically determines the length of the visual token sequence per image. The dynamic tokenization significantly reduces redundant information while preserving important visual cues, improving training efficiency and generation quality in tasks such as captioning, visual question answering, and image generation. DreamLLM [34], VL-GPT [35], MM-Interleaved [157], PUMA [158], and BLIP3-o [159] utilize OpenAI-CLIP encoder [22]. DreamLLM [34] introduces a lightweight linear projection to align CLIP embeddings with language tokens, while VL-GPT [35] employs a powerful casual transformer after OpenAI-CLIP vision encoder to effectively retain both semantic information and pixel details of the original image. Both MM-Interleaved [157] and PUMA [158] extract multi-granular image features via a CLIP tokenizer with simple ViT-Adapter or pooling operation to provide fine-grained feature fusion, thus supporting rich multimodal generation. BLIP3-o [159] reveals that flow matching loss is more effective than MSE loss, enabling more diverse image sampling and yielding better image quality. Mini-Gemini [160] introduces a visual token enhancement mechanism that requires dual semantic encoders. Specifically, it leverages a CLIP-pretrained ViT encoder [22] to obtain global visual tokens, while a LAION-pretrained ConvNeXt encoder provides dense local visual information. A cross-attention module is then employed to refine the global visual tokens by incorporating detailed visual cues from the dense encoder. These enhanced global tokens are subsequently combined with text tokens and processed by an LLM for joint vision-language understanding and generation. This design effectively bridges the semantic abstraction of CLIP features with the pixel-level precision of dense encoders. MetaMorph [161] employs SigLIP [152] to extract visual embeddings and introduces modality-specific adapters within a pretrained language model. These adapters are inserted throughout multiple transformer layers, allowing for deeper vision-language interaction compared to shallow projection approaches. ILLUME [162] adopt UNIT [154] as its vision encoder to provide a unified representation that balances semantic alignment and pixel-level fidelity. Unlike CLIP-like encoders that focus purely on contrastive objectives, UNIT [154] is jointly trained with both image reconstruction and contrastive alignment losses, producing tokens suitable for both vision-language understanding and image synthesis. Built on the powerful UNIT tokenizer, ILLUME effectively generates image tokens that retain both semantic and pixel-level information, which achieves better performance in multiple understanding and generation tasks, including captioning, VQA, text-to-image, and interleaved generation. Similarly, VILA-U [163] and UniTok [164] mimic UNIT [154] to introduce image-text contrastive learning to obtain a novel text-aligned vision tokenizer that balances semantic alignment and pixel-level fidelity. Recent work DualToken [165] use shallow-layer features of SigLIP for reconstruction and deep-layer features of SigLIP for semantic learning, thereby obtaining the texture and semantic visual features simultaneously. As a result, DualToken [165] achieves superior performance in both reconstruction and semantic tasks while demonstrating remarkable effectiveness in downstream MLLM understanding and generation tasks.
Across most of these models, causal attention masks are applied during MLLM training, and next-token prediction loss is used to optimize both text and vision token generation. For image generation, most of these models typically employ diffusion-based decoders, such as SD family [14, 166], IP-adapter [167], FLUX [16], and Lumina-Next [168], which are trained independently from the MLLM. During inference, the MLLM produces semantic-level visual tokens, which are then passed to the diffusion decoder for final image synthesis. This design choice—pairing semantic encoders with diffusion decoders—is motivated by the fact that semantic embeddings encode high-level conceptual information but lack the spatial density and low-level granularity required for direct pixel reconstruction. Diffusion models, with their iterative denoising mechanisms, are particularly well-suited for this setting: they are capable of progressively refining semantic representations into high-resolution, photorealistic images, even when the input tokens are sparse or abstract. In contrast, although few approaches (i.e., VILA-U [163] and UniTok [164]) adopt pixel-based decoders, their generated image quality is less competitive than the diffusion decoders. Thus, diffusion decoders provide a more robust and expressive decoding pathway for semantically compressed visual tokens, significantly improving text-image alignment, global coherence, and visual fidelity.
Current implementations of learnable query encoding can be broadly divided into two representative paradigms. The first is represented by SEED [169], which proposes a seed tokenizer that learns causal visual embeddings. Specifically, an input image is first encoded into dense token features via a BLIP-2 ViT encoder [53]. These features are then concatenated with a set of learnable query tokens and processed by a causal Q-Former to produce causal visual embeddings. This design is trained using both image-text contrastive learning and image reconstruction supervision, allowing the learned embeddings to simultaneously retain low-level visual detail and capture high-level semantic alignment with text. Building on this foundation, SEED-LLAMA [170] and SEED-X [171] enhance the model’s capacity by replacing the OPT backbone [172] with a stronger LLaMA2 model [2] and upgrading the decoder to UnCLIP-SD [14] or SDXL [166], leading to improved performance in both understanding and generation tasks. The second approach, introduced by MetaQueries [173], provides a simplified version of learnable query encoding. Here, image features are extracted via a frozen SigLIP encoder [152], which are then concatenated with learnable query tokens and directly passed through a frozen vision-language backbone such as LLaVA [136] or Qwen2.5-VL [10]. The output causal embeddings are used as conditioning inputs for a diffusion-based image decoder, enabling high-quality image generation. Because the backbone is kept frozen, the vision-language understanding capabilities remain consistent with the underlying pretrained models, offering a lightweight yet effective solution for multimodal generation. Nexus-Gen [174] and Ming-Lite-Uni [175] follow the MetaQueries paradigm, but with notable advancements to further enhance multimodal generation. Nexus-Gen [174] introduces a more powerful diffusion decoder, FLUX-1.dev, which significantly improves the generation quality. This approach allows the model to better capture the intricate details and high-fidelity features necessary for complex image generation tasks. On the other hand, Ming-Lite-Uni [175] takes a different route by introducing a highly capable MLLM model, M2-omini [176], for enhanced vision-language interaction. This model performs advanced vision-language conditioning to generate the conditioned image embeddings, ensuring a more semantically aligned representation. In addition, Ming-Lite-Uni fine-tunes its diffusion model by incorporating multi-scale learnable tokens, which facilitate improved semantic alignment across various visual scales. The multi-scale representation alignment mechanism enhances the model’s ability to generate detailed and contextually rich images from textual prompts, addressing challenges such as resolution mismatches and semantic inconsistencies. This innovative approach makes Ming-Lite-Uni a powerful tool for multimodal understanding and generation, pushing the boundaries of current methods in both flexibility and performance. To sum up, these learnable query-based designs share a common strength: they provide adaptive, compact, and semantically enriched representations that support both efficient image understanding and high-quality generation. By focusing on task-driven token extraction, such models offer a flexible and extensible alternative to traditional visual tokenizers, especially in unified multimodal frameworks.
Pseudo Hybrid Encoding. Representative works in this category include Janus [177], Janus-Pro [178], OmniMamba [179], and Unifluid [180]. As shown in Fig. 5 (b-4), these models adopt dual encoders—typically a semantic encoder (e.g., SigLIP) and a pixel encoder (e.g., VQGAN or VAE)—but use them in a task-specific manner. During training, the semantic encoder is enabled for vision-language understanding tasks, while the pixel encoder is activated for image generation tasks. Although the dual encoders are trained jointly with mixed understanding and generation data, the pixel encoder are not used at inference time. The motivation behind such a design is that mixed training with both types of data can enhance performance on both understanding and generation tasks. However, since only one encoder is active at a time, these models do not fully exploit the potential of hybrid encoding. In particular, they miss the opportunity to leverage semantic grounding in generation tasks and high-fidelity visual details in comprehension tasks. As such, these models typically employ pixel decoders to reconstruct images from latent codes.
Joint Hybrid Encoding. As shown in Fig. 5 (b-5), joint hybrid encoding methods integrate both semantic and pixel tokens into a single unified input for the language model or decoder, enabling simultaneous utilization of both representations. These models differ in their fusion strategies. MUSE-VL [181] and UniToken [182] concatenates the features from SigLIP and VQGAN along the channel dimension before passing them into the LLM. Tokenflow [183] incorporate dual encoders and codebooks with a shared mapping, enabling the joint optimization of high-level semantics and low-level pixel details. VARGPT [184], VARGPT-1.1 [185], and ILLUME+ [186] concatenate the semantic and pixel tokens along the sequence dimension, maintaining both token types in the LLM’s input. By integrating both semantic and detailed visual information, joint hybrid encoding enables more robust and expressive modeling capabilities for multimodal understanding and generation. These models support pixel decoders (e.g., VQGAN, Infinity [187], VAR-D30 [113]) as well as diffusion-based decoders (e.g., SDXL [166]), allowing them to generate images with improved semantic alignment and visual realism.
3.3 Fused Autoregressive and Diffusion Models
@@
BLIP3-o [159] b-2 Qwen2.5-VL OpenAI-CLIP Lumina-Next Causal 2025-05

### v4

@@ 2.2 Text-to-Image Model
FUDOKI [130] a DeepSeek-LLM SigLIP VQGAN VQGAN Bidirect. 2025-05
Muddit [131] a Meissonic (MM-DiT) VQGAN VQGAN Bidirect. 2025-05
@@
UniFork [153] b-2 Qwen2.5 SigLIP+RQ RQ-VAE Causal 2025-06
UniCode2 [154] b-2 Qwen2.5 SigLIP+RQ FLUX.1-dev / SD-1.5 Causal 2025-06
UniWorld [155] b-2 Qwen2.5-VL SigLIP2 DiT Bidrect. 2025-06
Pisces [156] b-2 LLaMA-3.1 SigLIP EVA-CLIP Diffusion Causal 2025-06
Tar [157] b-2 Qwen2.5 SigLIP2+VQ VQGAN / SANA Causal 2025-06
OmniGen2 [158] b-2 Qwen2.5-VL SigLIP OmniGen Causal 2025-06
@@
BLIP3-o [165] b-3 Qwen2.5-VL OpenAI-CLIP Learnable Query Lumina-Next Causal 2025-05
OpenUni [166] b-3 InternVL3 SigLIP Learnable Query SANA-1.5 Causal 2025-05
Ming-Omni [167] b-3 Ling QwenViT Learnable Query Multi-scale DiT Causal 2025-06
@@
Show-o2 [180] b-5 Qwen2.5 Wan-3DVAE + SigLIP Wan-3DVAE Wan-3DVAE Causal 2025-06
@@
3 Unified Multimodal Models for Understanding and Generation
3.1 Diffusion Models
Unlike Dual Diffusion [127], which combines discrete text diffusion with continuous image diffusion via Stable Diffusion [14], UniDisc [128] employs a fully discrete diffusion framework to train a Diffusion Transformer [199] from scratch. It tokenizes text using the LLaMA2 tokenizer [2] and converts images into discrete tokens with the MAGVIT-v2 encoder [200], allowing unification of both modalities in a discrete token space. These tokens undergo a discrete forward diffusion process, where structured noise is added simultaneously across modalities. In the reverse process, UniDisc progressively denoises the tokens to generate coherent sequences. The LLaMA2 and MAGVIT-v2 decoders then transform these sequences into high-quality text and images. By adopting a fully discrete approach, UniDisc enables simultaneous refinement of text and image tokens, enhancing inference efficiency and supporting versatile cross-modal conditioning.
In contrast to earlier discrete diffusion-based methods, FUDOKI [130] introduces a novel generative approach based on a discrete flow matching [201]. Under this framework, FUDOKI models a direct path between noise and data distributions by employing a kinetic-optimal, metric-induced probability trajectory. This design enables a continuous self-correction mechanism, which provides a clear advantage over the simple masking strategies used in earlier models. FUDOKI’s model architecture is based on Janus-1.5B [168]. However, it introduces essential modifications to support unified vision-language discrete flow modeling. One key change is the replacement of the standard causal mask with a full attention mask. This allows every token to attend to all others, thereby enhancing global contextual understanding. Although this modification removes the explicit causal structure, the model still supports next-token prediction by shifting its output logits by one position. Another important distinction is in the way FUDOKI handles time or corruption levels. Instead of relying on explicit time-step embeddings, as required in diffusion models, FUDOKI infers the corruption state directly from the input data. Following Janus-1.5B, FUDOKI decouples the processing paths for understanding and generation. A SigLIP encoder [202] is employed to capture high-level semantic features for image understanding, while a VQGAN-based tokenizer from LlamaGen [24] encodes the image into a sequence of low-level discrete tokens for image generation. At the output stage, the feature embeddings generated by the Janus-1.5B backbone are passed through modality-specific output heads to produce the final text and image outputs.
In a similar vein, Muddit [131] introduces a unified model for bidirectional generation using a purely discrete diffusion framework to handle text and images. Its architecture features a single Multimodal Diffusion Transformer (MM-DiT) with an architectural design similar to that of FLUX [203]. To leverage a strong image prior, the MM-DiT generator is initialized from Meissonic [204], a model extensively trained for high-resolution synthesis. Both modalities are quantized into a shared discrete space, where a pre-trained VQ-VAE [32] encodes images into codebook indices and a CLIP model [22] provides text token embeddings. During its unified training, Muddit employs a cosine scheduling strategy to mask tokens, and the single MM-DiT generator is trained to predict the clean tokens conditioned on the other modality. For output, a lightweight linear head decodes text tokens, while the VQ-VAE decoder reconstructs the image, allowing a single set of parameters to handle both text and image generation.
Despite these innovative approaches, significant challenges and limitations persist in the landscape of unified discrete diffusion models. A primary concern is inference efficiency. Although models like Mercury [207] and Gemini Diffusion [208] demonstrate potential for high-speed parallel token generation, most open-source discrete diffusion models still lag behind the practical inference speeds of their autoregressive counterparts. This discrepancy is primarily due to a lack of support for key-value cache and the degradation in output quality that occurs when decoding multiple tokens in parallel. The effectiveness of diffusion models is also hindered by training difficulties. Unlike autoregressive training, where every token provides a learning signal, discrete diffusion training offers only sparse supervision, as the loss is computed on a randomly selected subset of masked tokens, leading to inefficient use of the training corpus and high variance. Moreover, these models exhibit a length bias and struggle to generalize across different output lengths because they lack a built-in stopping mechanism like the end-of-sequence token found in autoregressive models. Additional development is also needed in architecture and supporting infrastructure. Architecturally, many existing models reuse designs originally created for autoregressive systems, an approach chosen for engineering simplicity that is not always suited to the diffusion process, which aims to capture joint data distributions in a way that is fundamentally different from the sequential nature of autoregressive models. On the infrastructure side, support for discrete diffusion models remains limited. Compared to the mature frameworks available for autoregressive models, they lack well-developed pipelines and robust open-source options. This gap hinders fair comparisons, slows research, and complicates real-world deployment. Addressing these interconnected challenges in inference, training, architecture, and infrastructure is essential to advance the capabilities and practical use of unified discrete diffusion models.
3.2 Auto-Regressive Models
2) Semantic Encoding. To overcome the semantic limitations inherent in pixel-based encoders, a growing body of work adopts semantic encoding, where image inputs are processed using pretrained text-aligned vision encoders such as OpenAI-CLIP [22], SigLIP [202], EVA-CLIP [36], or more recent unified tokenizers like UNIT [216], as shown in Fig. 5 (b-2). Some of these models leverage the multimodal features encoded by the multimodal autoregressive model as conditions for a diffusion model, enabling image generation with retained multimodal understanding capabilities, like OmniGen2 [158] that utilizes Qwen2.5-VL [10] as the multimodal model and enhanced OmniGen [217] as the image diffusion model. However, most of these models are trained on large-scale image-text pairs with contrastive or regression-based objectives, producing visual embeddings that align closely with language features in a shared semantic space. Such representations enable more effective cross-modal alignment and are particularly beneficial for multimodal understanding and generation.
Several representative models leverage different semantic encoders and architectural designs to support unified multimodal tasks. Emu [142], Emu2 [33], and LaViT [143] all employ EVA-CLIP [36] as their vision encoder. Notably, Emu [142] introduces the initial architecture combining a frozen EVA-CLIP encoder, a large language model, and a diffusion decoder to unify VQA, image captioning, and image generation. Emu2 [33] builds upon Emu [142] by proposing a simplified and scalable modeling framework for unified multimodal pretraining. It scales the MLLM model up to 37B parameters, significantly enhancing both understanding and generation capabilities. LaViT [143] introduces a dynamic visual tokenization mechanism built on top of EVA-CLIP. It employs a selector and merger module to adaptively select visual tokens from image embeddings based on content complexity. This process dynamically determines the length of the visual token sequence per image. The dynamic tokenization significantly reduces redundant information while preserving important visual cues, improving training efficiency and generation quality in tasks such as captioning, visual question answering, and image generation. DreamLLM [34], VL-GPT [35], and MM-Interleaved [144], and PUMA [147] utilize OpenAI-CLIP encoder [22]. DreamLLM [34] introduces a lightweight linear projection to align CLIP embeddings with language tokens, while VL-GPT [35] employs a powerful casual transformer after OpenAI-CLIP vision encoder to effectively retain both semantic information and pixel details of the original image. Both MM-Interleaved [144] and PUMA [147] extract multi-granular image features via a CLIP tokenizer with simple ViT-Adapter or pooling operation to provide fine-grained feature fusion, thus supporting rich multimodal generation. Mini-Gemini [145] introduces a visual token enhancement mechanism that requires dual semantic encoders. Specifically, it leverages a CLIP-pretrained ViT encoder [22] to obtain global visual tokens, while a LAION-pretrained ConvNeXt encoder provides dense local visual information. A cross-attention module is then employed to refine the global visual tokens by incorporating detailed visual cues from the dense encoder. These enhanced global tokens are subsequently combined with text tokens and processed by an LLM for joint vision-language understanding and generation. This design effectively bridges the semantic abstraction of CLIP features with the pixel-level precision of dense encoders. MetaMorph [148] employs SigLIP [202] to extract visual embeddings and introduces modality-specific adapters within a pretrained language model. These adapters are inserted throughout multiple transformer layers, allowing for deeper vision-language interaction compared to shallow projection approaches. ILLUME [149] adopt UNIT [216] as its vision encoder to provide a unified representation that balances semantic alignment and pixel-level fidelity. Unlike CLIP-like encoders that focus purely on contrastive objectives, UNIT [216] is jointly trained with both image reconstruction and contrastive alignment losses, producing tokens suitable for both vision-language understanding and image synthesis. Built on the powerful UNIT tokenizer, ILLUME effectively generates image tokens that retain both semantic and pixel-level information, which achieves better performance in multiple understanding and generation tasks, including captioning, VQA, text-to-image, and interleaved generation. Similarly, VILA-U [146] and UniTok [150] mimic UNIT [216] to introduce image-text contrastive learning to obtain a novel text-aligned vision tokenizer that balances semantic alignment and pixel-level fidelity. QLIP [151] addresses the potential conflict between reconstruction and text-image alignment tasks by implementing binary-spherical quantization. Tar [157] initiates the visual codebook by leveraging the vocabulary of LLMs and incorporates scale-adaptive pooling and decoding methodologies. This approach enables the model to adjust the length of the tokenizer according to the requirement: employing coarse-grained tokenizers for efficient generation and fine-grained tokenizers for comprehensive understanding. In generation tasks, Tar utilizes diffusion techniques to enhance the visual generation outcomes of AR models. UniFork [153] capitalizes on the text-aligned vision features of VILA-U. However, differentiating itself from the fully-shared parameters of understanding and generation MLLM, UniFork shares the parameters solely with these tasks at the shallow layer. At the deeper layer, these tasks are managed by distinct networks. This architecture successfully mediates the equilibrium between shared learning and task-specific specialization. UniCode2 [154] employs a cascaded codebook. In line with the method outlined in [218], it utilizes a substantial codebook derived from clustered SigLIP feature as the frozen foundational codebook, while introducing supplementary learnable codebooks to refine semantics specific to particular tasks. This separation enhances utilization and fosters robust learning. Recent work DualToken [152] uses shallow-layer features of SigLIP for reconstruction and deep-layer features of SigLIP for semantic learning, thereby obtaining the texture and semantic visual features simultaneously. As a result, DualToken [152] achieves superior performance in both reconstruction and semantic tasks while demonstrating remarkable effectiveness in downstream MLLM understanding and generation tasks.
Across most of these models, causal attention masks are applied during MLLM training, and next-token prediction loss is used to optimize both text and vision token generation. For image generation, most of these models typically employ diffusion-based decoders, such as SD family [14, 219], IP-adapter [220], FLUX [16], and Lumina-Next [221], which are trained independently from the MLLM. During inference, the MLLM produces semantic-level visual tokens, which are then passed to the diffusion decoder for final image synthesis. This design choice—pairing semantic encoders with diffusion decoders—is motivated by the fact that semantic embeddings encode high-level conceptual information but lack the spatial density and low-level granularity required for direct pixel reconstruction. Diffusion models, with their iterative denoising mechanisms, are particularly well-suited for this setting: they are capable of progressively refining semantic representations into high-resolution, photorealistic images, even when the input tokens are sparse or abstract. In contrast, although few approaches (i.e., VILA-U [146] and UniTok [150]) adopt pixel-based decoders, their generated image quality is less competitive than the diffusion decoders. Thus, diffusion decoders provide a more robust and expressive decoding pathway for semantically compressed visual tokens, significantly improving text-image alignment, global coherence, and visual fidelity. UniWorld [155] and Pisces [156] have endeavored to develop and expand such a solution. UniWorld directly utilizes the output features of pre-trained MLLM for visual comprehension as a high-level conditional signal, while employing SigLIP as low-level conditional signal to deliver comprehensive semantic visual control for DiT. Pisces employs EVA-CLIP as a condition for visual generation tasks and leverages diffusion to further enhance the model’s visual generation output. For various tasks, Pisces introduces tailored visual vector lengths and employs distinct MLPs to encode conditions. This approach increases the flexibility of model design while mitigating the inference cost compared to a single encoder configuration.
Current implementations of learnable query encoding can be broadly divided into two representative paradigms. The first is represented by SEED [159], which proposes a seed tokenizer that learns causal visual embeddings. Specifically, an input image is first encoded into dense token features via a BLIP-2 ViT encoder [53]. These features are then concatenated with a set of learnable query tokens and processed by a causal Q-Former to produce causal visual embeddings. This design is trained using both image-text contrastive learning and image reconstruction supervision, allowing the learned embeddings to simultaneously retain low-level visual detail and capture high-level semantic alignment with text. Building on this foundation, SEED-LLAMA [160] and SEED-X [161] enhance the model’s capacity by replacing the OPT backbone [222] with a stronger LLaMA2 model [2] and upgrading the decoder to UnCLIP-SD [14] or SDXL [219], leading to improved performance in both understanding and generation tasks. The second approach, introduced by MetaQueries [162], provides a simplified version of learnable query encoding. Here, image features are extracted via a frozen SigLIP encoder [202], which are then concatenated with learnable query tokens and directly passed through a frozen vision-language backbone such as LLaVA [209] or Qwen2.5-VL [10]. The output causal embeddings are used as conditioning inputs for a diffusion-based image decoder, enabling high-quality image generation. Because the backbone is kept frozen, the vision-language understanding capabilities remain consistent with the underlying pretrained models, offering a lightweight yet effective solution for multimodal generation. Open-Uni [166] refines the architecture of MetaQueries by utilizing solely learnable queries and a lightweight connector between a MLLM and a diffusion model, facilitating cohesive multimodal understanding and generation. Open-Uni demonstrates that the connetor between the MLLM visual understanding component and the diffusion-based visual generation component can be minimal in complexity, exemplified by a configuration comprising merely six Transformer layers. Nexus-Gen [163] and Ming-Lite-Uni [164] follow the MetaQueries paradigm, but with notable advancements to further enhance multimodal generation. Nexus-Gen [163] introduces a more powerful diffusion decoder, FLUX-1.dev, which significantly improves the generation quality. This approach allows the model to better capture the intricate details and high-fidelity features necessary for complex image generation tasks. On the other hand, Ming-Lite-Uni [164] takes a different route by introducing a highly capable MLLM model, M2-omini [193], for enhanced vision-language interaction. This model performs advanced vision-language conditioning to generate the conditioned image embeddings, ensuring a more semantically aligned representation. In addition, Ming-Lite-Uni fine-tunes its diffusion model by incorporating multi-scale learnable tokens, which facilitate improved semantic alignment across various visual scales. The multi-scale representation alignment mechanism enhances the model’s ability to generate detailed and contextually rich images from textual prompts, addressing challenges such as resolution mismatches and semantic inconsistencies. This innovative approach makes Ming-Lite-Uni a powerful tool for multimodal understanding and generation, pushing the boundaries of current methods in both flexibility and performance. Ming-Omni [167] adheres to the integrated MoE architecture, wherein modality-specific routing is facilitated through dedicated mechanisms tailored for each token, thereby enabling customized routing distributions. To address the multi-scale phenomenon inherent in visual generation [113], Ming-Omni employs multi-scale learnable queries, directed by an alignment strategy, to iteratively generate images progressing from coarse to fine detail. Furthermore, Ming-Omni integrates the audio modality and implements a dual-stage training strategy to mitigate the mutual influence between audio comprehension and generation tasks. The initial stage emphasizes comprehension capabilities, while the subsequent stage concentrates on enhancing generation quality. BLIP3-o [165] also employs learnable queries to bridge multimodal understanding and generation. However, it utilizes two diffusion models: one for learning CLIP embeddings and the other for using CLIP as a condition to generate images. It reveals that flow matching loss is more effective than MSE loss, enabling more diverse image sampling and yielding better image quality. To sum up, these learnable query-based designs share a common strength: they provide adaptive, compact, and semantically enriched representations that support both efficient image understanding and high-quality generation. By focusing on task-driven token extraction, such models offer a flexible and extensible alternative to traditional visual tokenizers, especially in unified multimodal frameworks.
Pseudo Hybrid Encoding. Representative works in this category include Janus [168], Janus-Pro [169], OmniMamba [170], Unifluid [171], and MindOmni [172]. As shown in Fig. 5 (b-4), these models adopt dual encoders—typically a semantic encoder (e.g., SigLIP) and a pixel encoder (e.g., VQGAN or VAE)—but use them in a task-specific manner. During training, the semantic encoder branch is enabled for vision-language understanding tasks, while the pixel encoder branch is activated for image generation tasks. Although the dual encoders are trained concurrently with combined understanding and generation datasets, the pixel encoder is not utilized during inference in understanding tasks and the semantic encoder is disabled for text-to-image generation. However, for image editing, Unifluid [171] uses the semantic encoder to encode the source image while MindOmni [172] utilizes both VAE and semantic encoder to encode the source image. The rationale behind this design choice is that mixed training with both types of data can enhance performance across understanding and generation tasks. However, since only one encoder is active at any given time, these models do not fully harness the advantages of hybrid encoding. Specifically, they miss the opportunity to employ semantic grounding in generation tasks and fail to utilize high-fidelity visual details in comprehension tasks. Consequently, these models typically engage pixel decoders to reconstruct images from latent codes.
Joint Hybrid Encoding. As shown in Fig. 5 (b-5), joint hybrid encoding methods integrate both semantic and pixel tokens into a single unified input for the language model or decoder, enabling simultaneous utilization of both representations. These models differ in their fusion strategies. MUSE-VL [173] and UniToken [179] concatenates the features from SigLIP and VQGAN along the channel dimension before passing them into the LLM. Tokenflow [174] incorporate dual encoders and codebooks with a shared mapping, enabling the joint optimization of high-level semantics and low-level pixel details. VARGPT [175], VARGPT-1.1 [177], and ILLUME+ [178] concatenate the semantic and pixel tokens along the sequence dimension, maintaining both token types in the LLM’s input. SemHiTok [176] introduces the Semantic Guided Hierarchical Codebook (SGHC), which perfectly inherits the semantic information of the semantic codebook while incorporating texture information to achieve pixel reconstruction. It is significant to observe that, contrary to other methods that directly employ distinct network branches for image processing, Show-o2 [180] utilizes separate network branches for the processing of latent features generated by 3DVAE [223], and uses the spatial-temporal fusion module to aggregate the outputs of different branches. This approach enables Show-o2 to capture both low-level and high-level visual information. However, such an operation might result in the loss of subtle semantic elements, owing to Show-o2’s use of 3D VAE for lossy compression of images or videos, potentially causing suboptimal handling of visual semantic details. By integrating both semantic and detailed visual information, joint hybrid encoding enables more robust and expressive modeling capabilities for multimodal understanding and generation. These models support pixel decoders (e.g., VQGAN, Infinity [224], VAR-D30 [113]) as well as diffusion-based decoders (e.g., SDXL [219]), allowing them to generate images with improved semantic alignment and visual realism.
3.3 Fused Autoregressive and Diffusion Models